OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 70 miles per hour (mph), and the other car is traveling 50 mph. How long will it take before the cars are 240 miles apart? 4 hrs 6 hrs 2 hrs 5 hrs

OpenStudy (anonymous):

is it 2? @some_someone

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

it is relativity if 2 cars run in opposite direction then their relative speed is the sum of 2's car speed :) so total speed is 70+50=120mph distance is 240miles so time= distance/speed=240/120=2hrs :)